Express each number as a product of its prime factors: 
156

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

156
By Taking the LCM of 156, we will get the product of its prime factor.
Hence, 156 = 2 × 2 × 13 × 3 × 1 = 2^2× 13 × 3

Thanks